# Paley-Wiener Theorems For Slice Regular Functions

### v1

@@ 3.2. Non-compact type Paley-Wiener theorem
Let $F:\mathbb{R}\to\mathbb{H}$ and $F\in L^{p}(\mathbb{R})$ for $1\leq p\leq 2$. Assume that there exists a quaternionic imaginary unit $I\in\mathbb{S}$ such that $\operatorname{supp}\mathscr{F}_{I}\left(F\right)\subset\left(-\infty,0\right]$. Then for every $J\in\mathbb{S}$, and $x>0,y\in\mathbb{R},$
is in $H^{p}\left(\mathbb{H}_{+}\right)$, where $\mathscr{F}_{E}\left(f\right)=\mathscr{F}_{I}\left(F\right)$.
@@
(3.12) $K\left(z\right)=\frac{1}{2\pi}\int_{-\infty}^{0}e^{zt}\mathrm{d}t,\ P\left(x,y\right)=\frac{|K\left(z\right)|^{2}}{K\left(2x\right)}.$
@@
The proof is similar to that given in [27]. To prove the necessity of Lemma 3.9, we need to show that, if $\Psi\left(x+Iy\right)\in H^{p}\left(\Pi_{+,I}\right)$, then $\operatorname{supp}\mathscr{F}_{I}\left(F\right)\subset\left(-\infty,0\right]$.
@@
The Proposition 2.17 implies that
@@
Now we turn to prove the sufficiency of Lemma 3.9. We have to show that if $F\in L^{p}(\mathbb{R})$ and $\operatorname{supp}\mathscr{F}_{I}\left(F\right)\subset\left(-\infty,0\right]$, then the relation (3.11) holds, and, as a consequence, $\Psi\left(x+Iy\right)\in H^{p}\left(\Pi_{+,I}\right)$.
@@
where $\chi_{\pm}\left(t\right)$ are characteristic functions on $[0,\infty)$ and $(-\infty,0]$ respectively. Since
@@
we have
$I\left(z,w\right)=$ $\int_{\mathbb{R}}\chi_{-}\left(t\right)\mathscr{F}_{I}\left(F\right)\left(t\right)\int_{\mathbb{R}}\chi_{-}\left(\xi\right)e^{z\xi}\chi_{+}\left(t-\xi\right)e^{-w\left(t-\xi\right)}\mathrm{d}\xi\mathrm{d}t$
$=$ $\int_{\mathbb{R}}\chi_{-}\left(t\right)\mathscr{F}_{I}\left(F\right)\left(t\right)\int_{\mathbb{R}}\chi_{-}\left(\xi\right)e^{z\xi}\chi_{-}\left(\xi-t\right)e^{w\left(\xi-t\right)}\mathrm{d}\xi\mathrm{d}t.$
@@
$I\left(z,w\right)=$ $\int_{\mathbb{R}}\chi_{-}\left(t\right)\mathscr{F}_{I}\left(F\right)\left(t\right)\int_{\mathbb{R}}\chi_{-}\left(s\right)e^{z\left(s+t\right)}e^{ws}\mathrm{d}s\mathrm{d}t$
$=$ $\int_{\mathbb{R}}\chi_{-}\left(t\right)e^{zt}\mathscr{F}_{I}\left(F\right)\left(t\right)\int_{\mathbb{R}}\chi_{-}\left(s\right)e^{\left(z+w\right)s}\mathrm{d}s\mathrm{d}t$
$=$ $\Psi\left(z\right)K\left(z+w\right).$
For $z=x+Iy\in\Pi_{+,I}$, we have $\overline{z}=x-Iy\in\Pi_{+,I}$, and
@@
where $P\left(x,y\right)$ is the Poisson kernel associated with the right half-plane $\Pi_{+,I}$. Due to (3.18), by $\lVert\Psi\rVert_{L^{p}(\mathbb{R})}=\lVert P*F\rVert_{L^{p}(\mathbb{R})}\leq\lVert P\rVert_{L^{1}(\mathbb{R})}\lVert F\rVert_{L^{p}(\mathbb{R})}$, we have
@@
So, we have $f\in H^{p}\left(\mathbb{H}_{+}\right)$. Moreover, if $f\in\mathscr{N}_{l}(\mathbb{H}_{+})\cap H^{p}(\mathbb{H}_{+})$, by Lemma 3.10, we have that for every $J\in\mathbb{S}$ and $x>0,y\in\mathbb{R}$,
@@ 4. Whittaker–Kotelnikov–Shannon Sampling Theorem
In the following we first define the Paley-Wiener space $PW_{A}^{2}\left(\mathbb{H}\right)$, $A>0$, where
@@
For all $\mathbf{q}=x+Iy\in\mathbb{H}$, the sinc function is defined by
@@
For all $I\in\mathbb{S}$ and $x,y\in\mathbb{R}$, we can get that
$\frac{1}{2}\left(\frac{\partial}{\partial x}+I\frac{\partial}{\partial y}\right)\mathrm{sinc}(x+Iy)=0$
and
$\mathrm{sinc}(x-Iy)=\overline{\mathrm{sinc}(x+Iy)}.$
Hence, we have $\mathrm{sinc}\in\mathscr{N}_{l}(\mathbb{H})$.
@@
Lemma 4.3.
For $M>0$, $p>1$ and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$,
$\sum_{k\in\mathbb{Z}}|\mathrm{sinc}\left(\mathbf{q}-k\right)|^{p}<p^{\prime}e^{pM\pi},\ \ \mathbf{q}\in\left\{\mathbf{q}=x+Jy\in\mathbb{H}:x\in\mathbb{R},|y|<M\right\}.$
Proof.
Let $|y|<M$, $\mathbf{q}=x+Jy$. Since the period of $\sum_{n\in\mathbb{Z}}\mathrm{sinc}\left(\mathbf{q}-n\right)$ is 1, we only need to consider $-\frac{1}{2}<x<\frac{1}{2}$. For such $x$, and $n\neq 0$, we have
$\lvert\mathrm{sinc}\left(\left(x+Jy\right)-n\right)\rvert$ $=\Big{|}\frac{\sin\pi\left(\left(x+Jy\right)-n\right)\cdot\left(\left(x-Jy\right)-n\right)}{\pi\lvert\left(x+Jy\right)-n\rvert^{2}}\Big{|}$
$=\frac{\lvert\sin\pi\left(\left(x+Jy\right)-n\right)\rvert}{\pi\lvert\left(x+Jy\right)-n\rvert}$
$=\frac{\Big{|}\frac{-J}{2}\left[e^{J\pi\left(\left(x+Jy\right)-n\right)}-e^{-J\pi\left(\left(x+Jy\right)-n\right)}\right]\Big{|}}{\pi\lvert\left(x+Jy\right)-n\rvert}$
$=\frac{\lvert e^{J\left(x+Jy\right)\pi}-e^{-J\left(x+Jy\right)\pi}\rvert}{2\pi\lvert\left(x+Jy\right)-n\rvert}$
$=\frac{\lvert e^{Jx\pi}e^{-y\pi}-e^{-Jx\pi}e^{y\pi}\rvert}{2\pi\lvert\left(x+Jy\right)-n\rvert}$
$\leq\frac{\lvert e^{Jx\pi}e^{-y\pi}\rvert+\lvert e^{-Jx\pi}e^{y\pi}\rvert}{2\pi\lvert\left(x+Jy\right)-n\rvert}$
$\leq\frac{e^{|y|\pi}}{\pi\lvert\left(x+Jy\right)-n\rvert}$
$\leq\frac{e^{M\pi}}{\pi\lvert x-n\rvert},$
By invoking the corresponding result in [24, Lemma 11.2], we have
$\sum_{k\in\mathbb{Z}}|\mathrm{sinc}\left(\left(x+Jy\right)-n\right)|^{p}$ $\leq e^{p|y|\pi}\left(1+\frac{1}{\pi^{p}}\sum_{k=1}^{\infty}\left\{\frac{1}{\left(n-x\right)^{p}}+\frac{1}{\left(n+x\right)^{p}}\right\}\right)$
$\leq p^{\prime}e^{pM\pi}.$
@@
Moverover, the convergence is absolute and uniform on strips of bounded width parallel to the real line, and on compact sets of $\mathbb{H}$.
@@
Combining (4), (4.3) and Lemma 4.3, we conclude that the sampling series is absolutely and uniformly convergent on strips of bounded width parallel to the real line. ∎
@@
Since $\frac{\sqrt{2\pi}}{2A}\sum_{k\in\mathbb{Z}}e^{Jkt}f\left(-\frac{\pi k}{A}\right)$ is the Fourier series of $\mathscr{F}_{J}\left(f|_{\mathbb{R}}\right)\in L^{2}\left(-A,A\right)$, we have that
$\frac{\sqrt{2\pi}}{2A}f\left(-\frac{\pi k}{A}\right)=\frac{1}{2A}\int_{-A}^{A}e^{-J\frac{\pi k}{A}t}\mathscr{F}_{J}\left(f|_{\mathbb{R}}\right)\left(t\right)\mathrm{d}t$
by the inversion Fourier formula. Hence the sampling series converges to $f$ in the $L^{2}(\mathbb{R})$-norm. ∎

### v2

@@ 3.2. Non-compact type Paley-Wiener theorem
Let $F:\mathbb{R}\to\mathbb{H}$ and $F\in L^{p}(\mathbb{R})$ for $1\leq p\leq 2$. Assume that there exists a quaternionic imaginary unit $I\in\mathbb{S}$ such that $\operatorname{supp}\mathscr{F}_{I}\left(F\right)\subset\left(-\infty,0\right]$. Then there exists a function $f\in H^{p}(\mathbb{H}_{+})$, whose non tangential boundary boundary limit at $Iy$ is $F(y)$, whence $\mathscr{F}_{E}\left(f\right)=\mathscr{F}_{I}\left(F\right)$ and
$f\left(x+Jy\right):=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{0}e^{\left(x+Jy\right)t}\mathscr{F}_{I}\left(F\right)\left(t\right)\mathrm{d}t,$
for all $x>0,y\in\mathbb{R}$, $J\in\mathbb{S}$.
@@
(3.9) $K\left(z\right)=\frac{1}{2\pi}\int_{-\infty}^{0}e^{z\xi}\mathrm{d}\xi,\quad P\left(x,y\right)=\frac{|K\left(z\right)|^{2}}{K\left(2x\right)}.$
@@
The proof is similar to that given in [29]. Let us first prove that, if $\Psi\left(x+Iy\right)\in H^{p}\left(\Pi_{+,I}\right)$, then $\operatorname{supp}\mathscr{F}_{I}\left(F\right)\subset\left(-\infty,0\right]$.
@@
Proposition 2.19 implies that
@@
Let us now prove that if $F\in L^{p}(\mathbb{R})$ and $\operatorname{supp}\mathscr{F}_{I}\left(F\right)\subset\left(-\infty,0\right]$, then Formula (3.8) holds, and, as a consequence, $\Psi\left(x+Iy\right)\in H^{p}\left(\Pi_{+,I}\right)$.
@@
where $\chi_{\pm}\left(t\right)$ are characteristic functions on $[0,\infty)$ and $(-\infty,0]$ respectively. Note that $\operatorname{supp}\mathscr{F}_{I}(F)\subset(-\infty,0]$, which means that $\mathscr{F}_{I}(F)(t)=\chi_{-}(t)\mathscr{F}_{I}(F)(t)$. Since
@@
and Lemma 3.7 we have
$I\left(z,w\right)=$ $\frac{1}{2\pi}\int_{\mathbb{R}}\chi_{-}\left(t\right)\mathscr{F}_{I}\left(F\right)\left(t\right)\int_{\mathbb{R}}\chi_{-}\left(\xi\right)e^{z\xi}\chi_{+}\left(t-\xi\right)e^{-w\left(t-\xi\right)}\mathrm{d}\xi\mathrm{d}t$
$=$ $\frac{1}{2\pi}\int_{\mathbb{R}}\chi_{-}\left(t\right)\mathscr{F}_{I}\left(F\right)\left(t\right)\int_{\mathbb{R}}\chi_{-}\left(\xi\right)e^{z\xi}\chi_{-}\left(\xi-t\right)e^{w\left(\xi-t\right)}\mathrm{d}\xi\mathrm{d}t.$
@@
$I\left(z,w\right)=$ $\frac{1}{2\pi}\int_{\mathbb{R}}\chi_{-}\left(t\right)\mathscr{F}_{I}\left(F\right)\left(t\right)\int_{\mathbb{R}}\chi_{-}\left(s\right)e^{z\left(s+t\right)}e^{ws}\mathrm{d}s\mathrm{d}t$
$=$ $\frac{1}{2\pi}\int_{\mathbb{R}}\chi_{-}\left(t\right)e^{zt}\mathscr{F}_{I}\left(F\right)\left(t\right)\int_{\mathbb{R}}\chi_{-}\left(s\right)e^{\left(z+w\right)s}\mathrm{d}s\mathrm{d}t.$
As a consequence, $I(z,w)=\Psi(z)K(z+w)$. For $z=x+Iy\in\Pi_{+,I}$, we have $\overline{z}=x-Iy\in\Pi_{+,I}$, and
@@
where $P\left(x,y\right)$ is the Poisson kernel associated with the right half-plane $\Pi_{+,I}$. Due to (3.15), by $\lVert\Psi(x+I\cdot)\rVert_{L^{p}(\mathbb{R})}=\lVert P*F\rVert_{L^{p}(\mathbb{R})}\leq\lVert P\rVert_{L^{1}(\mathbb{R})}\lVert F\rVert_{L^{p}(\mathbb{R})}$, we have
@@
So, we have $f\in H^{p}\left(\mathbb{H}_{+}\right)$. Moreover, since $f\in\mathscr{N}_{l}(\mathbb{H}_{+})\cap H^{p}(\mathbb{H}_{+})$, by Lemma 3.9, we have that for every $J\in\mathbb{S}$ and $x>0,y\in\mathbb{R}$,
@@ 4. Whittaker–Kotelnikov–Shannon Sampling Theorem
The classical complex Paley-Wiener space, the sampling theorem and the sinc function play important roles in Fourier analysis, numerical analysis and signal analysis (see [26, 31, 40]). So it is the motivation for us to study the sampling theorem for slice regular functions. In the following we first define the Paley-Wiener space $PW_{A}^{2}\left(\mathbb{H}\right)$, $A>0$, where
@@
For all $\mathbf{q}\in\mathbb{H}$, the sinc function is defined by
$\mathrm{sinc}(\mathbf{q})=\sum_{k=0}^{\infty}\mathbf{q}^{2k}\frac{(-\pi^{2})^{k}}{(2k+1)!}=(\pi\mathbf{q})^{-1}\sin(\pi\mathbf{q})=\sin(\pi\mathbf{q})(\pi\mathbf{q})^{-1}.$
It follows from Proposition 2.4 that $\mathrm{sinc}(\mathbf{q})\in\mathscr{N}_{l}(\mathbb{H})$.
Remark 4.2.
For all $\mathbf{q}=x+Iy\in\mathbb{H}$, the integral expression of sinc is given by
@@
Remark 4.4.
Here, the sinc function in Definition 4.1 has a property analogous to the complex sinc function, as shown below: For $M>0$, $p>1$ and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$, we have
$\sum_{k\in\mathbb{Z}}|\mathrm{sinc}\left(\mathbf{q}-k\right)|^{p}<p^{\prime}e^{pM\pi}$
for all $\mathbf{q}\in\Omega_{M}:=\left\{\mathbf{q}=x+Jy\in\mathbb{H}:x\in\mathbb{R},|y|<M\right\}.$ The proof is similar to that given in [5, Lemma 2.2]. It is useful to prove Theorem 4.5.
@@
Moverover, the convergence is absolute and uniform on $\Omega_{M}=\{\mathbf{q}=x+Jy\in\mathbb{H}:x\in\mathbb{R},|y|<M\}$ for $M<\infty$, and on compact sets of $\mathbb{H}$.
@@
Combining (4), (4.3) and Remark 4.4, we conclude that the sampling series is absolutely and uniformly convergent on $\Omega_{M}=\left\{\mathbf{q}=x+Jy\in\mathbb{H}:x\in\mathbb{R},|y|<M\right\}$ for $M<\infty$. ∎
@@
Using Proposition 2.11, we have $f=f_{0}+f_{1}\tilde{I}+f_{2}\tilde{J}+f_{3}\tilde{I}\tilde{J}$ where $f_{m}\in\mathscr{N}_{l},m=0,1,2,3$ and $\tilde{I},\tilde{J}$ are mutually orthogonal quaternionic imaginary units. It follows that
$\begin{split}&\Big{\|}\mathscr{F}_{J}\left(f|_{\mathbb{R}}\right)\left(t\right)-\frac{\sqrt{2\pi}}{2A}\sum_{|k|\leq n}e^{J\frac{\pi k}{A}t}f\left(-\frac{\pi k}{A}\right)\Big{\|}_{L^{2}\left(-A,A\right)}\\ =&\Big{\|}\mathscr{F}_{J}(f_{0}|_{\mathbb{R}})(t)-\frac{\sqrt{2\pi}}{2A}\sum_{|k|\leq n}e^{J\frac{\pi k}{A}t}f_{0}\left(-\frac{\pi k}{A}\right)\\ &+\mathscr{F}_{J}(f_{1}|_{\mathbb{R}})(t)\tilde{I}-\frac{\sqrt{2\pi}}{2A}\sum_{|k|\leq n}e^{J\frac{\pi k}{A}t}f_{1}\left(-\frac{\pi k}{A}\right)\tilde{I}\\ &+\mathscr{F}_{J}(f_{2}|_{\mathbb{R}})(t)\tilde{I}-\frac{\sqrt{2\pi}}{2A}\sum_{|k|\leq n}e^{J\frac{\pi k}{A}t}f_{2}\left(-\frac{\pi k}{A}\right)\tilde{J}\\ &+\mathscr{F}_{J}(f_{3}|_{\mathbb{R}})(t)\tilde{I}\tilde{J}-\frac{\sqrt{2\pi}}{2A}\sum_{|k|\leq n}e^{J\frac{\pi k}{A}t}f_{3}\left(-\frac{\pi k}{A}\right)\tilde{I}\tilde{J}\Big{\|}_{L^{2}\left(-A,A\right)}\\ \leq&\sum_{m=0}^{3}\Big{\|}\mathscr{F}_{J}\left(f_{m}|_{\mathbb{R}}\right)\left(t\right)-\frac{\sqrt{2\pi}}{2A}\sum_{|k|\leq n}e^{J\frac{\pi k}{A}t}f_{m}\left(-\frac{\pi k}{A}\right)\Big{\|}_{L^{2}\left(-A,A\right)}.\end{split}$
Since $f\in PW_{A}^{2}(\mathbb{H})$, we have $f_{m}\in PW_{A}(\mathbb{H})$. Note that
$\frac{\sqrt{2\pi}}{2A}f_{m}\left(-\frac{\pi k}{A}\right)=\frac{1}{2A}\int_{-A}^{A}e^{-J\frac{\pi k}{A}t}\mathscr{F}_{J}\left(f_{m}|_{\mathbb{R}}\right)\left(t\right)\mathrm{d}t.$
Thus, $\frac{\sqrt{2\pi}}{2A}\sum_{k\in\mathbb{Z}}e^{J\frac{\pi k}{A}t}f_{m}\left(-\frac{\pi k}{A}\right)$ is the Fourier series of $\mathscr{F}_{J}\left(f_{m}|_{\mathbb{R}}\right)\in L^{2}\left(-A,A\right)$. It is easy to see that
$\lim_{n\to\infty}\Big{\|}\mathscr{F}_{J}\left(f_{m}|_{\mathbb{R}}\right)\left(t\right)-\frac{\sqrt{2\pi}}{2A}\sum_{|k|\leq n}e^{J\frac{\pi k}{A}t}f_{m}\left(-\frac{\pi k}{A}\right)\Big{\|}_{L^{2}\left(-A,A\right)}=0.$
Hence the sampling series converges to $f$ in the $L^{2}(\mathbb{R})$-norm. ∎